# An unidentified Fermi source emitting radio bursts in the Galactic bulge

### v1

@@ 5.2.3 Young Pulsars
Another class of -ray emitting neutron stars is young pulsars with periods between from fifty to a few hundred milliseconds. These can be prominent giant-pulse emitters. Spectral indices of giant pulses from Crab are typically –1.443.5 (Karuppusamy et al., 2010). Giant pulses are very narrow pulses with widths ranging from nanoseconds to microseconds. Some studies have shown that young pulsars are more than 70% linearly polarized. Circular polarization has also been detected in single pulses of young pulsars (Kramer et al., 2002). No obvious orthogonal jumps were reported in most and most of them also show flat PA swings (Johnston & Weisberg, 2006). Young pulsars are capable of emitting very bright single pulses too (Large et al., 1968). The polarization fraction, non-detection of PA jumps, brightness of single pulses, and -ray luminosities are consistent with the source being a young pulsar.
@@ 5.2.4 Rotating Radio Transients
Rotating Radio Transients (RRATs) are sparsely emitting, characterized by their detection more prominently in single-pulse rather than Fourier-based searches (McLaughlin et al., 2006; Burke-Spolaor, 2013). They have spin periods ranging 45ms – 7.7s and surface magnetic fields in the range – G (Abhishek et al., 2022). One RRAT has been previously detected in X-rays (Rea et al., 2009), but none has been associated with -rays yet. RRATs show diverse polarization properties, with some single pulses having 100% linear polarization fraction and a significant amount of circular polarization (Hsu et al., 2023). Some RRATs also show discontinuous PA jumps as expected from orthogonal polarization modes (Caleb et al., 2019), which we do not see in our source.
@@ 5.2.5 Binary star system
While the Pan-STARRS star association (§5.1) was inconclusive, it is worth considering the implications if the source were in a binary. The M-dwarf nature of the star will make the system a low-mass binary. -ray emission from low mass binaries suggests that the neutron star is a recycled MSP. But we disfavor an MSP as the source because of the reasons outlined in §5.2.2. However, GeV emissions have been detected from high-mass binaries like PSR B1259–63 and PSR J2032+4127. The GeV emission is suggested to arise from a combination of Bremsstrahlung and inverse Compton emission from unshocked and weakly shocked electrons of the pulsar wind (Chang et al., 2021). The isotropic -ray luminosity during the periastron of PSR B1259–63 is approximately 10 times higher than the observed Fermi luminosity of J1818–1531 (Abdo et al., 2011). A hidden high-mass companion star would be required to explain the GeV emission of J1818–1531. Eclipses by the companion star could also help explain the intermittency of pulses seen from this object, and a line-of-sight through a variable plasma medium (as during orbit) could provide the DM and scattering variations we observed.
@@ 6 Conclusion
Here we reported the discovery of a source of radio bursts discovered by the realfast system at VLA during commensal operation on the VLA. Using realfast we were also able to localize the bursts to subarcsecond precision. Follow-up observations with GBT detected more bursts but did not yield any clear periodicity. The bursts exhibit a generally downward-trending radio luminosity at higher radio frequencies. The brightness temperature and duration of the bursts suggest a compact source with a coherent emission mechanism. All the bursts at 1.4 GHz shows significant scattering, which is consistent with what is expected from the DM. However, the scattering timescale changed significantly between the first and second epochs, suggesting a dynamically evolving scattering screen. The bursts shows an average RM of 816 rad m. All bursts showed high linear polarization and significant circular polarization was detected in one burst. We compared the field of this source with multi-wavelength catalogs, finding that the realfast bursts appear confidently linked to the Fermi source, and are spatially coincident with a Pan-STARRS star (however, the data provide only inconclusive confidence in their relationship). This also revealed an XMM-Newton source 1.9 away from the realfast position, nearby but not consistent with the burst source’ position. Overall, it is expected that the burst source is of neutron star origin based on its clear similarities to the variety of radio-emitting pulsar phenomena.
Overall, the polarization fraction and radio spectrum is consistent with most magnetars, single pulses from MSPs and young pulsars, and RRATs. The flat or gradually changing behavior of polarization angle and the radio energies are similar to the population of young pulsars and magnetars. However, the high -ray luminosities favor a young pulsar or a binary origin of J1818–1531. The source could be a magnetar, however this will be the first detection of a magnetar in GeV energies. Targeted X-ray follow-up is required to strengthen the magnetar origin and an optical follow-up would best explore the source as a binary, however further monitoring of pulsed radio behaviors may improve the chances at detecting a periodicity in this source and providing identification via pulsar timing.

### v2

@@ 5.2.3 Young Pulsars
Another class of -ray emitting neutron stars is young pulsars with periods between from fifty to a few hundred milliseconds. These can be prominent giant-pulse emitters. Giant pulses are very narrow pulses with widths ranging from nanoseconds to microseconds. Some studies have shown that young pulsars are more than 70% linearly polarized. Circular polarization has also been detected in single pulses of young pulsars (Kramer et al., 2002). No obvious orthogonal jumps were reported in most and most of them also show flat PA swings (Johnston & Weisberg, 2006). Young pulsars are capable of emitting very bright single pulses too (Large et al., 1968). The polarization fraction, non-detection of PA jumps, brightness of single pulses, and -ray luminosities are consistent with the source being a young pulsar.
@@ 5.2.4 Rotating Radio Transients
Rotating Radio Transients (RRATs) are sparsely emitting, characterized by their detection more prominently in single-pulse rather than Fourier-based searches (McLaughlin et al., 2006; Burke-Spolaor, 2013). They have spin periods ranging 45ms – 7.7s and surface magnetic fields in the range – G (Abhishek et al., 2022). One RRAT has been previously detected in X-rays (Rea et al., 2009), but none has been associated with -rays yet. RRATs show diverse polarization properties, with some single pulses having 100% linear polarization fraction and a significant amount of circular polarization (Hsu et al., 2023). Some RRATs also show discontinuous PA jumps as expected from orthogonal polarization modes (Caleb et al., 2019), which we do not see in our source. However, in Fig. 6 the bursts from J1818–1531 occupy the same spaces as RRATs and if this is indeed an RRAT, it would be the first one to be discovered to have a ray counterpart.
@@ 5.2.5 Binary star system
While directly associating with the star described in §5.1 proved unlikely, it is worth considering the implications if the source were in a binary. If there is somehow (e.g. due to unrecognized localization error issues) and association between the realfast bursts and the star discussed in §5.1.1, the M-dwarf nature of the star will make the system a low-mass binary. -ray emission from low mass binaries suggests that the neutron star is a recycled MSP. But we disfavor an MSP as the source because of the reasons outlined in §5.2.2. However, GeV emissions have been detected from high-mass binaries like PSR B1259–63 and PSR J2032+4127. The GeV emission is suggested to arise from a combination of Bremmsstrahlung and inverse Compton emission from unshocked and weakly shocked electrons of the pulsar wind (Chang et al., 2021). The isotropic -ray luminosity during the periastron of PSR B1259–63 is consistent with the observed Fermi luminosity of J1818–1531 (Abdo et al., 2011). A hidden high-mass companion star would be required to explain the GeV emission of J1818–1531. Eclipses by the companion star or absorption of the radio emission from the pulsar by the outflows from the star (Chen et al., 2021) could also help explain the intermittency of pulses seen from this object, and a line-of-sight through a variable plasma medium (as during orbit) could provide the DM and scattering variations we observed.
@@ 6 Conclusion
Here we reported the discovery of a source of radio bursts discovered by the realfast system at VLA during commensal operation on the VLA. Using realfast we were also able to localize the bursts to subarcsecond precision. Follow-up observations with GBT detected more bursts but did not yield any clear periodicity. The brightness temperature and duration of the bursts suggest a compact source with a coherent emission mechanism. All the bursts at 1.4 GHz shows significant scattering, which is consistent with what is expected from the DM. However, the scattering timescale changed significantly between the first and second epochs, suggesting a dynamically evolving scattering screen. The bursts shows an average RM of 816 rad m-2. All bursts showed high linear polarization and significant circular polarization was detected in one burst. We compared the field of this source with multi-wavelength catalogs, finding that the realfast bursts appear confidently linked to the Fermi source, and are spatially coincident with a PAN-STARRS star (however, the data provide only inconclusive confidence in their relationship). This also revealed an XMM-Newton source 1.9′′ away from the realfast position, nearby but not consistent with the burst source’ position. Overall, it is expected that the burst source is of neutron star origin based on its clear similarities to the variety of radio-emitting pulsar phenomena.
Overall, the polarization property is consistent with most magnetars, single pulses from MSPs and young pulsars, and RRATs. The flat or gradually changing behavior of polarization angle and the radio energies are similar to the population of young pulsars and magnetars. However, the high -ray luminosities favor a young pulsar or a binary origin of J1818–1531. The source could be a magnetar, however this will be the first detection of a magnetar in GeV energies. Targeted X-ray follow-up is required to strengthen the magnetar origin and an optical follow-up would best explore the source as a binary, however further monitoring of pulsed radio behaviors may improve the chances at detecting a periodicity in this source and providing identification via pulsar timing.